Ben is planning to fence his rectangular garden. The area of the garden is 50 square feet, and the length of the garden is twice the width. How many feet of fencing will he need?
Answer by richwmiller(17219)

x*2x=50
3x=50
x=50/3
2x=100/3
2*2x+2x=f
6x=f
6*50/3=100 feet of fence